OpenStudy (anonymous):

A light ray in water is incident on a boundary with air at an angle of 22°. The refractive index of water is 1.33 and that of air is 1.00. What is the angle of refraction in air? 15° 22° 30° 45°

OpenStudy (anonymous):

you must use Snell's Law! \[\frac{ \sin \theta_1 }{ \sin \theta_2 } = \frac{ n_2 }{ n_1 }\]